# PagPassGPT: Pattern Guided Password Guessing via Generative Pretrained Transformer

### v1

@@ Abstract
Amidst the surge in deep learning-based password guessing models, challenges of generating high-quality passwords and reducing duplicate passwords persist. To address these challenges, we present PagPassGPT, a password guessing model constructed on Generative Pretrained Transformer (GPT). It can perform pattern guided guessing by incorporating pattern structure information as background knowledge, resulting in a significant increase in the hit rate. Furthermore, we propose D&C-GEN to reduce the repeat rate of generated passwords, which adopts the concept of a divide-and-conquer approach. The primary task of guessing passwords is recursively divided into non-overlapping subtasks. Each subtask inherits the knowledge from the parent task and predicts succeeding tokens. In comparison to the state-of-the-art model, our proposed scheme exhibits the capability to correctly guess 12% more passwords while producing 25% fewer duplicates.
@@ I Introduction
Florêncio et al. [3] investigated various user accounts and observed that the majority of accounts are not very sensitive. According to their research, users might opt to create a new account rather than spend 10 minutes recovering a lost one. Rather than being a target, users are more likely to face threats from trawling attacks.
@@ I-B2 D&C-GEN
  • We investigate the issue of trawling attacks and unveil the shortcomings of existing deep learning-based password guessing schemes. Additionally, we introduce PagPassGPT, which addresses these weaknesses by properly integrating deep learning-based models with password patterns. Furthermore, in an effort to reduce the occurrence of duplicate passwords, we propose D&C-GEN algorithm, which adopts a divide-and-conquer approach for the task of password guessing.
@@ II-A1 Trawling Attack
Trawling Attack [4] is one of the earliest attacks and has drawn substantial attention. In a trawling attack, the attacker does not specifically target an individual user but rather focuses on a broad group of users. The attacker builds password guessing models by modeling real leaked passwords [34, 35], and then uses the models to generate a large number of passwords to hit the new users’ real passwords. The attacker does not care about which user is under attack. The attacker’s objective is to maximize the hit rate while minimizing the number of guesses. Hence, password guessing models should prioritize generating passwords with higher probabilities of being used.
@@ III-B1 Training Process
As shown in the left part of Fig. 3, during the training process, each input password undergoes preprocessing followed by tokenization, both implemented within a specialized component called the tokenizer. As shown in the left part of Fig. 4, in the phase of training preprocessing, the tokenizer of PagPassGPT applies PCFG (detailed in Section II-C) to extract the password pattern, “L4N3S1”, from the input password, “Pass123\$”. “L4N3S1” stands for a password consisting of four letters followed by three numbers and one special character. Then, it utilizes the extracted pattern to concatenate with the password, forming a rule in the format as below,
@@
where <BOS> represents the beginning of the sequence, <SEP> denotes the separator, and <EOS> stands for the ending of the sequence. In the phase of tokenization, as shown in Fig. 5, the tokenizer serves two functions: encoding and decoding. During encoding, the tokenizer takes the former rule as input and produces tokenized indexes, while during decoding, it reverses the process, mapping tokenized indexes back to the rule. In particular, during the encoding phase, the tokenizer initially splits the input content into segments, considering each segment as a token. Each token is then mapped to an index based on a vocabulary, ensuring that every token has a unique index. For example, as shown in Fig. 5, the tokenizer splits the input rule into segments and each of them is mapped to an index according to the applied vocabulary, forming the index list $[0,25,14,40,1,66,77,95,95,42,43,44,106,2]$. The applied vocabulary consists of three categories of tokens: 94 visible ASCII character tokens, excluding the space character; 5 special tokens (<BOS>, <SEP>, <EOS>, <UNK>, and <PAD>); and 36 pattern tokens (e.g., L12, S12, and N12), totaling 136 tokens. <UNK> is a token used to represent an out-of-vocabulary token, and <PAD> is the padding token utilized to pad the indexes list.
@@ III-C D&C-GEN
D&C-GEN is proposed to reduce the repeat rate of generated passwords, inspired by the concept of the divide and conquer approach. In this section, we detail D&C-GEN algorithm in three parts. The first part is to illustrate its implementation and the second part is to analyze its effectiveness. Finally, we demonstrate its optimization.
@@ III-C1 Design
[…]
@@ IV-C Pattern Guided Guessing Test
The experiment of pattern guided guessing test is designed in five steps. The initial step involves computing the probability distribution of extracted patterns from passwords in the test set, and categorizing them based on different numbers of segments. Each segment represents a format of organizing characters.
@@
where $NH$ denotes the number of hits, ${TC}_{s}^{test}$ and ${TC}_{P}^{test}$ represents the total number of passwords from the test set conforming to a certain category with $s$ segments and conforming to a specific pattern $P$ respectively.
As shown in Fig. 8, with an increasing number of pattern segments, PagPassGPT consistently outperforms PassGPT. For instance, when the number of segments is $1$, the distinction between PagPassGPT and PassGPT is less pronounced. As the number of segments increases to $5$, the gap reaches its peak, with the hit rates ($HR_{s}$) of 13.00% for PassGPT and 40.54% for PagPassGPT. When the number of segments exceeds $9$, the hit rate of PassGPT approaches zero. However, PagPassGPT continues to demonstrate its utility. We further illustrate the details of the hit rate of each pattern in Fig. 9. For the convenience of presentation, we only show the top 5 patterns of each category from the first segment to the sixth segments. PagPassGPT demonstrates a higher hit rate ($HR_{P}$) for almost all patterns compared to PassGPT. Particularly, PagPassGPT is capable of guessing passwords with challenging patterns, while PassGPT fails to make any correct guesses.
@@
In order to illustrate the disparity in pattern guided guessing between PassGPT and PagPassGPT, we randomly select ten passwords generated by each model, adhering to the “L5N2” and “L5S1N2” patterns, as outlined in Table III. From the table, it is evident that passwords generated by PassGPT tend to exhibit word truncation, particularly when English words are involved. For example, in the password “polic#10”, the word “police” lacks the letter “e” because PassGPT must insert a special character in its subsequent token to adhere to the pattern requirement when the generation process reaches “e”. In contrast, PagPassGPT rarely encounters such issues since it considers not only the pattern requirement but also the model prediction.
@@ IV-E Cross-Site Attack Test
In order to evaluate the generality of the proposed model, we conduct a cross-site attack test. We first train the most recent PassGPT and the proposed PagPassGPT on Rockyou and LinkedIn independently. Then we evaluate them by testing the hit rates of $10^{8}$ passwords on other datasets and the result of hit rates is shown in Table VI. PassGAN, VAEPass, and PassFlow are excluded from the comparison since their hit rates show a significant gap with PassGPT and PagPassGPT at $10^{8}$ guesses in the trawling test. Specifically, their hit rates are less than 10% while the hit rate of both PassGPT and PagPassGPT are over 20% as shown in Table IV.

### v2

@@ Abstract
Amidst the surge in deep learning-based password guessing models, challenges of generating high-quality passwords and reducing duplicate passwords persist. To address these challenges, we present PagPassGPT, a password guessing model constructed on a Generative Pretrained Transformer (GPT). It can perform pattern guided guessing by incorporating pattern structure information as background knowledge, resulting in a significant increase in the hit rate. Furthermore, we propose D&C-GEN to reduce the repeat rate of generated passwords, which adopts the concept of a divide-and-conquer approach. The primary task of guessing passwords is recursively divided into non-overlapping subtasks. Each subtask inherits the knowledge from the parent task and predicts succeeding tokens. In comparison to the state-of-the-art model, our proposed scheme exhibits the capability to correctly guess 12% more passwords while producing 25% fewer duplicates.
@@ I Introduction
Florêncio et al. [3] investigated various user accounts and observed that the majority of accounts are not essential. According to their research, users might opt to create a new account rather than spend 10 minutes recovering a lost one. Rather than being a target, users are more likely to face threats from trawling attacks.
@@ I-B2 D&C-GEN
  • We investigate the issue of trawling attacks and unveil the shortcomings of existing deep learning-based password guessing schemes. Additionally, we introduce PagPassGPT, which addresses these weaknesses by properly integrating deep learning-based models with password patterns. Furthermore, in an effort to reduce the occurrence of duplicate passwords, we propose the D&C-GEN algorithm, which adopts a divide-and-conquer approach for the task of password guessing.
@@ II-A1 Trawling Attack
Trawling Attack [4] is one of the earliest attacks that has drawn substantial attention. In a trawling attack, the attacker does not specifically target an individual user but rather focuses on a broad group of users. The attacker builds password guessing models by modeling real leaked passwords [34, 35], and then uses the models to generate a large number of passwords to hit the new users’ real passwords. The attacker does not care about which user is under attack. The attacker’s objective is to maximize the hit rate while minimizing the number of guesses. Hence, password guessing models should prioritize generating passwords with higher probabilities of being used.
@@ III-B1 Training Process
As shown in the left part of Fig. 3, during the training process, each input password undergoes preprocessing followed by tokenization, both implemented within a specialized component called the tokenizer. As shown in the left part of Fig. 4, in the phase of training preprocessing, the tokenizer of PagPassGPT applies PCFG (detailed in Section II-C) to extract the password pattern, “L4N3S1”, from the input password, “Pass123\$”. “L4N3S1” stands for a password consisting of four letters followed by three numbers and one special character. Then, it utilizes the extracted pattern to concatenate with the password, forming a rule in the format below,
@@
where <BOS> represents the beginning of the sequence, <SEP> denotes the separator, and <EOS> stands for the ending of the sequence. In the phase of tokenization, as shown in Fig. 5, the tokenizer serves two functions: encoding and decoding. During encoding, the tokenizer takes the former rule as input and produces tokenized indexes, while during decoding, it reverses the process, mapping tokenized indexes back to the rule. In particular, during the encoding phase, the tokenizer initially splits the input content into segments, considering each segment as a token. Each token is then mapped to an index based on a vocabulary, ensuring that every token has a unique index. For example, as shown in Fig. 5, the tokenizer splits the input rule into segments, and each of them is mapped to an index according to the applied vocabulary, forming the index list $[0,25,14,40,1,66,77,95,95,42,43,44,106,2]$. The applied vocabulary consists of three categories of tokens: 94 visible ASCII character tokens, excluding the space character; 5 special tokens (<BOS>, <SEP>, <EOS>, <UNK>, and <PAD>); and 36 pattern tokens (e.g., L12, S12, and N12), totaling 136 tokens. <UNK> is a token used to represent an out-of-vocabulary token, and <PAD> is the padding token utilized to pad the indexes list.
@@ III-C D&C-GEN
D&C-GEN is proposed to reduce the repeat rate of generated passwords, inspired by the concept of the divide and conquer approach. In this section, we detail the D&C-GEN algorithm in three parts. The first part is to illustrate its implementation and the second part is to analyze its effectiveness. Finally, we demonstrate its optimization.
@@ III-C1 Design
The above description is detailed in Algorithm 1. It takes the total number $N$ of password guessing attempts, the threshold $T$ of dividing a task, and a set $S_{p}$ of patterns and their probabilities. For each pattern $P_{i}$, we first compute the number $N_{P_{i}}$ of passwords to be generated through the total number $N$ of attempts multiplying the probability $Pr(P_{i})$ of the pattern. After that, the comparison between the number of passwords to be generated and the threshold is conducted. If the number of passwords to be generated is smaller than the threshold, the task is directly executed and outputs the passwords to a set $R$. Otherwise, the pattern $P_{i}$ and the number $N_{P_{i}}$ of passwords to be generated conforming to the pattern $P_{i}$ is added into a list $L_{P_{i}}$. For each element ($\textit{Pref}_{i}$, $n_{i}$) of $L_{P_{i}}$, We first pop it from the list and then compare its number $n_{i}$ of passwords to be generated with the threshold $T$. If it is smaller than the threshold, the task is executed and outputs the passwords to $R$. Otherwise, the model is executed to get the probability of following $c$ tokens with the current prefix Pref. $c$ is the number of candidate tokens conforming to the current pattern requirement. In our setup, the variable $c$ is assigned different values: $52$ for a letter, $10$ for a number, and $32$ for a special character, depending on the type of the next token. After that, for each new token, the number $n_{j}$ of passwords to be generated is calculated and the current prefix is concatenated with the new token to form the new prefix $\textit{Pref}_{j}$. Subsequently, the number $n_{j}$ of passwords to be generated and the new prefix $\textit{Pref}_{j}$ are added into the list $L_{P_{i}}$. Finally, the algorithm outputs the set $R$. For clarity, we’ve included an example in Fig. 7 to illustrate the process of D&C-GEN.
@@ IV-C Pattern Guided Guessing Test
The experiment of pattern guided guessing test is designed in five steps. The initial step involves computing the probability distribution of extracted patterns from passwords in the test set and categorizing them based on different numbers of segments. Each segment represents a format for organizing characters.
@@
where $NH$ denotes the number of hits, ${TC}_{s}^{test}$ and ${TC}_{P}^{test}$ represent the total number of passwords from the test set conforming to a certain category with $s$ segments and conforming to a specific pattern $P$ respectively.
As shown in Fig. 8, with an increasing number of pattern segments, PagPassGPT consistently outperforms PassGPT. For instance, when the number of segments is $1$, the distinction between PagPassGPT and PassGPT is less pronounced. As the number of segments increases to $5$, the gap reaches its peak, with the hit rates ($HR_{s}$) of 13.00% for PassGPT and 40.54% for PagPassGPT. When the number of segments exceeds $9$, the hit rate of PassGPT approaches zero. However, PagPassGPT continues to demonstrate its utility. We further illustrate the details of the hit rate of each pattern in Fig. 9. For the convenience of presentation, we only show the top 5 patterns of each category from the first segment to the sixth segment. PagPassGPT demonstrates a higher hit rate ($HR_{P}$) for almost all patterns compared to PassGPT. Particularly, PagPassGPT is capable of guessing passwords with challenging patterns, while PassGPT fails to make any correct guesses.
@@
To illustrate the disparity in pattern guided guessing between PassGPT and PagPassGPT, we randomly select ten passwords generated by each model, adhering to the “L5N2” and “L5S1N2” patterns, as outlined in Table III. From the table, it is evident that passwords generated by PassGPT tend to exhibit word truncation, particularly when English words are involved. For example, in the password “polic#10”, the word “police” lacks the letter “e” because PassGPT must insert a special character in its subsequent token to adhere to the pattern requirement when the generation process reaches “e”. In contrast, PagPassGPT rarely encounters such issues since it considers not only the pattern requirement but also the model prediction.
@@ IV-E Cross-Site Attack Test
To evaluate the generality of the proposed model, we conduct a cross-site attack test. We first train the most recent PassGPT and the proposed PagPassGPT on Rockyou and LinkedIn independently. Then we evaluate them by testing the hit rates of $10^{8}$ passwords on other datasets and the result of hit rates is shown in Table VI. PassGAN, VAEPass, and PassFlow are excluded from the comparison since their hit rates show a significant gap with PassGPT and PagPassGPT at $10^{8}$ guesses in the trawling test. Specifically, their hit rates are less than 10% while the hit rates of both PassGPT and PagPassGPT are over 20% as shown in Table IV.